The amount of phosphorus present in a sample at a given time is given by the following equation:

Qt = Q0e-0.03t

What percent of the initial amount of phosphorus is left after 25 days of decay?
1
Expert's answer
2020-12-22T04:34:04-0500

"Q_t = Q_0e^{-0.03t} \\

\frac{Q_t}{Q_0} = e^{-0.03 \times 25} \\

\frac{Q_t}{Q_0} = e^{-0.75} \\

\frac{Q_t}{Q_0} = e^{-0.75} \\

\frac{Q_t}{Q_0} = 0.472"

Proportion:

Q0 – 100%

Qt – x

"x = \frac{Q_t \times 100}{Q_0} \\

= 0.472 \times 100 = 47.2" %

Answer: 47.2 %